A circular wire loop of radius R carries a total charge q distributed uniformly over its length. A small length x(very very less than R) of the wire is cut off. Find the electric field at the centre due to the remaining wire.

A  circular wire loop of radius R carries a total charge q distributed uniformly over its length. A small length x(very very less than R) of the wire is cut off. Find the electric field at the centre due to the remaining wire.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Total charge given for loop = q.
Charge per unit length = q/2*pi*R
for dl element,
dq = qdl/2*pi*a.
Net electricf field at center due to that small length and reaining is:
0 = E + kdq/r^2
thus just put dq value now.